# MOND as a transformation between non-inertial reference frames via Sciama’s interpretation of Mach’s Principle

### v2

@@ Abstract
Milgrom’s Modified Newtonian Dynamics (MOND) correction to Newtonian gravity or inertia is shown to be equivalent to a more fundamental transformation between a non-inertial local reference frame and the fixed background of the observable universe, complying with Mach’s principle. Both Newton’s gravitational constant and Milgrom’s MOND acceleration parameter or scale constant are substituted for two varying, measurable, physical, and cosmological parameters under the justification of Sciama’s interpretation of Mach’s principle: causally connected mass and size of the universe. This Machian interpretation, scale-invariant and free from fundamental constants and free parameters with the exception of the speed of light as the speed of causality and gravity, is based on relative field intensities of the small and large scale of the universe. It respects the Machian effect from Sciama’s model, by which, in absence of a background, rotational speed is undefined up to the speed of light. The Machian MOND approximation is a necessary feature of a phenomenological nonlinear theory of modified inertia or nonlinear theory of modified gravity which incorporates Mach’s principle in agreement with galaxy rotation curves.
@@ 1 Introduction
1.1 Modified Newtonian Dynamics
@@
The solution to the problem of the perihelion precession of Mercury first involved various failed attempts proposing the unobserved existence of a new small planet Vulcan, an asteroid belt, or a gas cloud orbiting the Sun. However, the resolution to the problem came as a modification to the laws of gravity by General Relativity (GR). Similarly, the observation of rotational velocities in disks of galaxies higher than predicted by Newtonian gravity, as depicted in galaxy rotation curves, has driven further research into modifications to the laws of gravity. The discrepancies between observed luminance to gravitational mass ratio in galaxies and their missing Keplerian falloff in their velocity curves, together with faster radial velocities of galaxies in galaxy clusters than Newtonian predictions, among other evidence, are referred to as the “dark matter” problem [1, 2, 3]. Modifying gravity instead of proposing the existence of physical cold dark matter is motivated by the Renzo’s rule, the Tully-Fisher law, and the cuspy halo problem. These recent modifications to the classical laws of gravity are called Modified (Orbital) Newtonian Dynamics (MOND) hypotheses (or MilgrOmiaN Dynamics), the first and main MOND theory being Milgrom’s MOND [4, 5, 6], which can be thought of as a modification to inertia through Newton’s second law of motion, or to the inverse square law of gravity.
@@
The reason for the anomalous motion of Mercury was that Newtonian gravity is not a good enough approximation for strong gravitational fields, such as near the Sun, where nonlinear GR effects become significant. Milgrom’s MOND explains the observed rotation curves of galaxies by imposing a correction based on two observational relationships: The asymptotic velocities in rotation curves are proportional to the fourth root of the mass (Tully-Fisher law) and the mass discrepancies are always observed below a particular acceleration scale. The simplest correction that satisfies these observational conditions results in a true acceleration $g=GM/\mu\left({g}/{a_{0}}\right)r^{2}$, achieving on purpose a constant asymptotic rotational speed $v=\sqrt[4]{GMa_{0}}$. The connection between the classical Newtonian regime and Milgrom’s MOND low acceleration regime (at which the inverse-square law of gravity changes to a linear law) is determined by an “interpolating function” $\mu\left(g/a_{0}\right)=1/(1+\left(a_{0}/g\right)^{n})^{(1/n)}$ (with $n=1$ for the simple and $n=2$ for the standard interpolating function) based on the “true acceleration” $g$ and a constant of acceleration $a_{0}$ (or gravitational field intensity, also referred to as the Hubble acceleration or the acceleration scale constant), which is a free parameter adjusted by data fitting and is in principle considered a fundamental and universal constant in Milgrom’s MOND. The interpolating function must satisfy the conditions $\mu(x)\rightarrow 1$ for $x\ll 1$ and $\mu(x)\rightarrow x$ for $x\gg 1$. A fixed-distance or fixed-mass (independent of the system under study) modification alone cannot account for the observed rotation velocities due to the Tully-Fisher law, i.e., it is not possible to simply set a constant of distance or a constant of mass to modify the laws of gravity to fit all rotation curves of galaxies (because some galaxies exhibit the effect of dark matter at small radius while others at large radius, and same happens with different galaxy masses). But by combining both mass and distance in Newtonian acceleration or gravitational field intensity, Milgrom’s MOND achieves this limit where the effect of dark matter in most galaxies becomes relevant. Still, Milgrom’s MOND, being insufficient to explain all dark matter related phenomena such as galaxy cluster dynamics, [7, 8], being non-relativistic, and not satisfying conservation laws, is an effective theory or approximation to a more fundamental modification to the theory of gravity.
@@
The acceleration constant in Milgrom’s MOND $a_{0}\approx 1.2\cdot 10^{-10}\ m/s^{2}$, was already related to the cosmological scale through the Hubble constant $H_{0}$ (or the cosmological constant, respectively) $a_{0}\sim cH_{0}\sim c^{2}\sqrt{\Lambda}$ originally by Milgrom through the Hubble parameter and then through the cosmological constant [9], stating that this coincidence could point to a basic theory underlying MOND’s phenomenology. Milgrom stated that “an attractive possibility is that MOND results as a non-relativistic, small-scale expression of a fundamental theory by which inertia is a vestige of the interaction of a body with “the rest of the Universe”, in the spirit of Mach’s principle.[10], although he did not further develop his theory in relation to Mach’s principle.
@@
Various theories of modified gravity are based on Milgrom’s MOND acceleration constant: John Moffat’s MOG scalar–tensor–vector gravity approach to galaxy rotation curves is based on two parameters, which are set to fit galactic Milgrom’s MOND’s acceleration constant [11], Erik’s Verlinde’s entropic gravity attempts to propose an underlying framework for Milgrom’s MOND’s acceleration constant [12], and Deur’s self-interacting gravitons model correction to Newtonian gravity contains a physical constant counterpart to Milgrom’s MOND’s acceleration constant in the form of $a=\Lambda/a_{0}=\sqrt{\Lambda}/c^{2}$ [13]. Others, such as quantized inertia, rely on another acceleration parameter in the form of $a=c^{2}/R_{u}$ where $R_{u}$ is the radius of the cosmological comoving horizon [14].
1.2 Mach’s Principle
@@
Newtonian gravity and classical mechanics are founded on the principles of absolute space, absolute time, and absolute motion (including acceleration). These were originally opposed by Leibniz (and others, such as Berkeley in his 1721 De Motu, and Huygens) arguing that as observers, we can only epistemically access relative notions of space, time, and motion. Newton’s justification of absolute space and motion is portraited in his rotating bucket of water experiment at the beginning of his 1687 masterpiece Philosophiæ naturalis principia mathematica, which can be summarized as follows: A bucket of water is set spinning around its axis and, at first, the walls of the bucket rotate relative to the stationary water while the surface of the water remains flat as prior to the spinning. After the water starts to rotate as well, it rises towards the walls and its shape is no longer flat when the spinning bucket and the water are at rest relative to each other. Interactions between water and walls (due to their relative velocities) cannot explain the shape of the water, and Newton assumed that this experiment could be used to measure rotation with respect to an absolute space, “without relation to anything external, which remains always similar and immovable”.
@@
The definition of absolute space, time, and motion always suffers from circular reasoning. In Newtonian mechanics, absolute space is defined as the frame absent of inertial forces (also referred to as fictitious forces, which come from our choice of reference frame that is rotating in absolute space), but the absence of inertial forces is used to prove and identify the existence of absolute space in Newton’s rotating bucket of water experiment. According to Newton’s first law of motion, an object moves inertially if it is free from outside influences, but the fact that it is free from outside influences is inferred only by observing that it moves inertially. In Special Relativity, light travels at the same speed in all inertial frames and absolute acceleration is defined relative to inertial frames, but inertial frames are defined as frames absent of acceleration (accelerometers always measure acceleration with respect to a reference frame of calibration, which could be non-inertial). If a body is measured to have acceleration in one inertial frame, all other inertial frames will agree that it is accelerating.
@@
Leibniz’s relativistic space, time and motion ideas were further developed by Ernst Mach in his 1883 The Science of Mechanics (which inspired Einstein to develop GR, who originally coined the term “Mach’s principle”), in which he criticizes Newton’s conclusion of his bucket of water experiment stating that “No one is competent to say how the experiment would turn out if the sides of the vessel increased in thickness and mass till they were ultimately several leagues thick.[15]. According to Mach, the relative rotation of the water with respect to the bucket produces no noticeable centrifugal forces, and such forces are instead produced by its relative rotation with respect to Earth and the other celestial bodies (in this way, the relative motion between the several leagues thick bucket and the water could produce noticeable centrifugal forces in the water, and perhaps their non-relative motion could reduce these forces when both rotate with respect to the rest of the universe). In contrast to Newton, who attempted to explain the physical effects of inertia through a sort of resistance to motion within an unobservable absolute space with no physical properties, Mach conceived inertia as an interaction that required other external bodies to manifest. Newton did not suspect that the change in the water’s shape could be due to rotation relative to the rest of the universe because his action at a distance gravitational force from his infinite and homogeneous universe acting on the water cancels out according to Newton’s shell theorem and it is independent of velocity or acceleration. But no local phenomena can ever be isolated from the rest of the universe, which certainly reaches the water even if the sum of these forces is zero. According to Mach, and in opposition to Newton’s conclusions, a spinning bucket of water in an empty universe would not change the water’s shape (one could not detect relative motion in an empty universe, it would be undefinable), and if the rest of the universe was set spinning while the bucket was at rest, the water surface would curve, as both the spinning of the bucket or the spinning of the universe are indistinguishable systems without an absolute space.
@@
According to Mach, and known as Mach’s principle, the inertia of a body is not an independent and intrinsic property of matter (unlike in Newtonian mechanics and GR, where a particle in an empty universe has inertial properties), but rather the result of the action of the universe as a whole. Mach suggested that the fixed background distribution of matter in the universe must exert inertial forces on a local accelerating body. In this way, the reference system with respect to which the universe is at rest or in uniform and rectilinear motion is a true inertial reference system. In Ernst Mach words, “I have remained to the present day the only one who insists upon referring the law of inertia to the Earth, and in the case of motions of great spatial and temporal extent to the fixed stars.[15]. Hence, inertial frames should be defined with respect to this rest frame, and local physical laws must be determined by the large-scale structure of the universe.
@@
A frame linearly accelerated relative to an inertial frame in the Minkowskian spacetime of Special Relativity is locally identical to a frame at rest in a gravitational field. Einstein’s equivalence principle assumes that inertial and gravitational masses are equal, and the metric tensor in GR determines the inertial mass of a body. Gravity is the only long distance force that cannot be screened. Moreover, there always exists a reference frame in which inertial forces vanish, just as for the case of gravitational forces in free fall. These equivalences indicate that both inertial and gravitational forces are of the same nature. Following Mach, in any two-body interaction, the influence of all other matter inside their causal spacetime should be taken into account. Inertia is thus a form of gravitational induction, appearing when a body is accelerated with respect to the rest of the universe, subjected to retarded action. It seems that to describe, for example, the Earth’s motion around the Sun, only local masses and distances are required, but the Machian perspective implies that the universe’s action is already taken into account in the Newtonian laws. The only way this can be true is through the Newtonian gravitational constant, which, together with the choice of an inertial absolute space frame, both constitute the two arbitrarily choices of Newton.
@@
In Newtonian gravity, the gravitational constant $G$ is the only fundamental constant of the theory, and it is known with far less precision than any other fundamental dimensional constant in physics. The uncertainty in its measurement, together with its problematic in Quantum Field Theory in Planck’s energy (which is used as a cutoff for the energy density of the vacuum and the theoretical quantum corrections estimation of the Higg’s mass, leading to the vacuum catastrophe and hierarchy problems) has led to questioning its constant nature repeatedly over the last century, see [17] for reference. GR carries this constant (a varying $G$ would imply a violation of the strong equivalence principle) with the addition of the speed of light, which can be derived from vacuum permittivity and permeability constants. Therefore, it is reasonable to question whether the gravitational constant is also a derived parameter.
@@
The first historically suggestion of gravity arising necessarily as a consequence of the relativity of inertia came from Hans Reissner [18]. Schrödinger soon identified in 1925 [19] that GR did not fully implement Mach’s principle, and proposed a relationship between the gravitational potential of distant masses and the speed of light: “This remarkable relationship states that the (negative) potential of all masses at the point of observation, calculated with the gravitational constant valid at that observation point, must be equal to half the square of the speed of light.”. The conclusion was reached by imposing that the kinetic energy had an origin in a potential-like interaction (such as all other forms of energy, which have an origin in an interaction) and was relational, so that $K=mv^{2}/2$ based on absolute mass and velocity can be expressed through the gravitational Newtonian potential $Gm/r$ so that $K=(Gm_{i}/r_{ij})(m_{j}v_{ij}^{2}/2c^{2})$ with the necessity of introducing $c^{2}$. The term $Gm_{i}/r_{ij}c^{2}=GM_{u}/c^{2}R_{u}\sim 1$ is considered missing in the original formulation, with $M_{u}$ being the sum of all masses and $R_{u}$ the distance between the moving particle $m_{j}$ and $M_{u}$. This relationship was later credited to Reissner by Schrödinger himself.
@@
Schrödinger’s formulation appeared repeatedly in several works after him [20, 21, 22, 23, 24], but it was popularized by Dennis Sciama, who developed a modified cosmological vector potential theory of inertia in which the inertial law arises as a side effect of gravity [25]. He proposed that local inertial forces result from the gravitational induction of the universe, so that the dynamics of a rotating body would be affected by the large-scale distribution of mass of the universe. The induced inertia decreased with $1/r$ with $r$ being the distance to the inertial body, so that the action of global matter dominates over the action of local matter, and it is not significantly modified by the acceleration of local matter, leading to the illusion that inertia depends only on the body itself. He postulated that “In the rest-frame of any body the total gravitational field at the body arising from all the other matter in the universe is zero”. The gravitational constant at any point was determined by the total gravitational potential of the distribution of matter of the universe, taking the form of $G\sim c^{2}R_{u}/M_{u}=c^{2}/\Phi_{u}$ with $M_{u}$ and $R_{u}$ the mass and radius of the observable universe, causally connected by the speed of light $c$ in the past light-cone of the point considered to evaluate $G$, for the equivalence of inertial and gravitational masses to hold. In this way, local phenomena were strongly coupled to global properties of the universe. Sciama’s model predicted a non-infinite and expanding universe, since the potential of an infinite non-expanding eternal universe would also be infinite. Sciama’s relationship was also derived through an alternative approach based on the velocity of waves in a medium of a certain pressure and mass density, in which the rest energy of a given gravitational mass is its gravitational potential energy due to the distribution of masses of the universe [26].
@@
By pure dimensional analysis, Reissner’s and Schrödinger’s equation (henceforth Sciama’s relationship) is the only possible derivation of the units of measurement of the gravitational constant through a mass, a distance, and the speed of light. If Sciama’s relationship is derived from a more fundamental theory, it can be expected to also contain integers and mathematical constants, but these will be omitted for simplicity, since it does not significantly affect the resulting orders of magnitude of the comparison between $M_{u}$ and $R_{u}$.
@@
Sciama’s relationship with today’s measurements of the current cosmological model is satisfied only in terms of orders of magnitude, because the Hubble tension impedes a precise calculation. It yields the orders of magnitude of the observed gravitational constant when considering not the baryonic and dark matter content for $M_{u}$, but the total energy content (since all forms of energy must gravitate according to GR), which roughly corresponds to the critical density of the universe as measured to be almost flat. However, the critical density is calculated with the assumption of a constant gravitational constant itself and considering dark matter, which could be eliminated through a correction to the gravitational constant, as will be proposed in the next section. Thus, an honest calculation must be done with direct observations only by the total luminosity of the universe and its average mass-to-light ratio, which yields a value $10^{52}kg<M_{u}<10^{53}kg$, and the observable radius of the universe, which can be estimated by direct observations of the Hubble parameter (subjected to the Hubble tension), which yields around $R_{u}\sim 10^{26}m$.
@@
Few authors have attempted to develop a MOND for galaxy rotation curves based on Mach’s principle, or relate Milgrom’s MOND to Sciama’s relationship and Mach’s principle [27, 28, 29, 30, 31, 32, 33]. The most significant points in the literature relating these topics are summarized onwards.

### v3

@@ Abstract
Milgrom’s Modified Newtonian Dynamics (MOND) correction to Newtonian gravity or inertia is shown to be equivalent to a more fundamental transformation between a non-inertial local reference frame and the fixed background of the observable universe, complying with Mach’s principle. Both Newton’s gravitational constant and Milgrom’s MOND acceleration parameter or scale constant are substituted for two varying, measurable, physical, and cosmological parameters under the justification of Sciama’s interpretation of Mach’s principle: causally connected mass and size of the universe. This Machian interpretation of MOND, mass, length and time scale invariant at all regimes and free from fundamental constants and free parameters with the exception of the speed of light as the speed of causality and gravity, is based on relative field intensities of the small and large scale of the universe. It respects the Machian effect from Sciama’s modified inertia model, by which, in absence of a background, rotational speed is undefined up to the speed of light. The Machian MOND interpretation is a necessary feature that a phenomenological non-linear theory of modified inertia or non-linear theory of modified gravity which incorporates Mach’s principle in agreement with galaxy rotation curves should reduce to as an approximation.
@@ 1 Introduction
1.1 Modified Newtonian Dynamics MOND
@@
The solution to the problem of the perihelion precession of Mercury first involved various failed attempts proposing the unobserved existence of a new small planet Vulcan, an asteroid belt, or a gas cloud orbiting the Sun. However, the resolution to the problem came as a modification to the laws of gravity by General Relativity (GR). Similarly, the observation of rotational velocities in disks of galaxies higher than predicted by Newtonian gravity, as depicted in galaxy rotation curves, has driven further research into modifications to the laws of gravity and mechanics. The discrepancies between observed luminance to gravitational mass ratio in galaxies and their missing Keplerian falloff in their velocity curves, together with faster radial velocities of galaxies in galaxy clusters than Newtonian predictions, among other evidence, are referred to as the ‘dark matter’ problem [1, 2, 3]. Modifying gravity or inertia instead of proposing the existence of physical cold dark matter is motivated by the Tully-Fisher law, the Renzo’s rule, and the cuspy halo problem. These recent modifications to the classical law of gravity and/or laws of mechanics are called Modified (or Mondified) Newtonian Dynamics (MOND) hypotheses (or MilgrOmiaN Dynamics), the first and main MOND theory being Milgrom’s MOND [4, 5, 6], which can be thought of as a modification to inertia through Newton’s second law of motion, or to the inverse square law of gravity. Almost all full-fledged versions of MOND are just modified gravity and not modified inertia. As modified inertia, MOND modifies the kinetic particle Lagrangian term $mv^{2}/2$ or the ‘inertia term’ $mg$, and measured kinematic accelerations are not simply related to the gravitational forces per unit mass.
@@
The reason for the anomalous motion of Mercury was that Newtonian gravity is not a good enough approximation for strong gravitational fields, such as near the Sun, where non-linear GR effects become significant. Milgrom’s MOND generally explains the observed rotation curves of galaxies by imposing a correction based on two observational relationships: the asymptotic velocities in rotation curves are proportional to the fourth root of the mass (Tully-Fisher law), and the mass discrepancies are always observed below a particular acceleration scale. The simplest correction that satisfies these observational conditions results in a constant asymptotic rotational velocity $v=\sqrt[4]{GMa_{0}}$ with the need to introduce a constant of acceleration $a_{0}$. The connection between the classical Newtonian regime where $v=\sqrt{GM/r}$ and Milgrom’s deep MOND or low acceleration regime (at which, as modified gravity, the inverse-square law of gravity changes to a linear law) is determined by an ‘interpolating function’ $\mu\left(g/a_{0}\right)=1/(1+\left(a_{0}/g\right)^{n})^{(1/n)}$ (with $n=1$ for the simple and $n=2$ for the standard interpolating function) based on $g$ as the Newtonian acceleration in modified gravity or the ‘true acceleration’ in modified inertia and a constant of acceleration $a_{0}$ (also referred to as the Hubble acceleration or the acceleration scale constant), which is a free parameter adjusted by data fitting and is in principle considered a fundamental and universal constant in Milgrom’s MOND. The interpolating function must satisfy the conditions $\mu(x)\rightarrow 1$ for $x\ll 1$ and $\mu(x)\rightarrow x$ for $x\gg 1$. A fixed-distance or fixed-mass (independent of the system under study) modification alone cannot account for the observed rotation velocities due to the Tully-Fisher law, i.e., it is not possible to simply set a constant of distance or a constant of mass to modify the laws of gravity to fit all rotation curves of galaxies (because some galaxies exhibit the effect of dark matter at small radius while others at large radius, and same happens with different galaxy masses). But by combining both mass and distance in Newtonian acceleration or gravitational field intensity, Milgrom’s MOND achieves this limit where the effect of dark matter in galaxies is generally observed. Still, Milgrom’s MOND, being insufficient to explain all dark matter related phenomena such as galaxy cluster dynamics, [7, 8], being non-relativistic, and not explaining phenomenologically the origin of the interpolating function, is an effective theory or approximation to a more fundamental modification to gravitation or inertia.
@@
The acceleration constant in Milgrom’s MOND $a_{0}\sim 10^{-10}\ m/s^{2}$, was already related to the cosmological scale through the Hubble constant $H_{0}$ (or the cosmological constant $\Lambda$) $a_{0}\sim cH_{0}\sim c^{2}\sqrt{\Lambda}$ originally by Milgrom [9], stating that this coincidence could point to a basic theory underlying MOND’s phenomenology. In particular, he stated that “an attractive possibility is that MOND results as a non-relativistic, small-scale expression of a fundamental theory by which inertia is a vestige of the interaction of a body with ‘the rest of the Universe’, in the spirit of Mach’s principle.[10], although he did not further develop his theory in relation to Mach’s principle.
@@
Various theories of modified gravity or inertia are based on Milgrom’s MOND acceleration constant: John Moffat’s MOG scalar–tensor–vector gravity approach to galaxy rotation curves is based on two parameters, which are set to fit galactic Milgrom’s MOND’s acceleration constant [11], Erik’s Verlinde’s entropic gravity attempts to propose an underlying framework for Milgrom’s MOND’s acceleration constant [12], and Deur’s self-interacting gravitons model correction to Newtonian gravity contains a physical constant counterpart to Milgrom’s MOND’s acceleration constant in the form of $a=\Lambda/a_{0}=\sqrt{\Lambda}/c^{2}$ [13]. Others, such as quantized inertia, rely on another acceleration parameter in the form of $a=c^{2}/R_{u}$ where $R_{u}$ is the radius of the cosmological comoving horizon [14].
1.2 Mach’s Principle and modified inertia
@@
Newtonian gravity and classical mechanics are founded on the principles of absolute space, absolute time, and absolute motion (including acceleration). These were originally opposed by Leibniz (and others, such as Berkeley in his 1721 De Motu, and Huygens) arguing that as observers, we can only epistemically access relative notions of space, time, and motion. Newton’s justification of absolute space and motion is portraited in his rotating bucket of water experiment at the beginning of his 1687 masterpiece Philosophiæ naturalis principia mathematica, which can be summarized as follows: a bucket with water is set spinning around its axis and, at first, the walls of the bucket rotate relative to the stationary water while the surface of the water remains flat as prior to the spinning. After the water starts to rotate as well, it rises towards the walls and its shape is no longer flat when the spinning bucket and the water are at rest relative to each other. Interactions between water and walls (due to their relative velocities) cannot explain the shape of the water, and Newton assumed that this experiment could be used to measure rotation with respect to an absolute space, “without relation to anything external, which remains always similar and immovable”.
@@
The definition of absolute space, time, and motion suffer from circular reasoning. In Newtonian mechanics, absolute space is defined as the frame absent of inertial forces (also referred to as fictitious forces, which come from our choice of reference frame that is rotating in absolute space), but the absence of inertial forces is used to prove and identify the existence of absolute space in Newton’s rotating bucket of water experiment. According to Newton’s first law of motion, an object moves inertially if it is free from outside influences, but the fact that it is free from outside influences is inferred only by observing that it moves inertially. In Special Relativity (SR), light travels at the same speed in all inertial frames and absolute acceleration is defined relative to inertial frames, but inertial frames are defined as frames absent of acceleration (accelerometers always measure acceleration with respect to a reference frame of calibration, that must be assumed to be inertial). If a body is measured to have acceleration in one inertial frame, all other inertial frames will agree that it is accelerating.
@@
Leibniz’s relative space, time and motion ideas were further developed by Ernst Mach in his 1883 The Science of Mechanics (which inspired Einstein to develop GR, who originally coined the term ‘Mach’s principle’), in which he criticized Newton’s conclusion of his bucket of water experiment stating that “No one is competent to say how the experiment would turn out if the sides of the vessel increased in thickness and mass till they were ultimately several leagues thick.[15]. According to Mach, the relative rotation of the water with respect to the bucket produces no noticeable centrifugal forces, and such forces are instead produced by its relative rotation with respect to Earth and the other celestial bodies (in this way, the relative motion between the several leagues thick bucket and the water could produce noticeable centrifugal forces in the water, and perhaps their non-relative motion could reduce these forces when both rotate with respect to the rest of the universe). In contrast to Newton, who attempted to explain the physical effects of inertia through a sort of resistance to motion within an unobservable absolute space with no physical properties, Mach conceived inertia as an interaction that required other external bodies to manifest. Newton did not suspect that the change in the water’s shape could be due to rotation relative to the rest of the universe because his action at a distance gravitational force from his infinite and homogeneous universe acting on the water cancels out according to Newton’s shell theorem and it is independent of velocity or acceleration. But no local phenomena can ever be isolated from the rest of the universe, and its effect certainly reaches the water even if the sum of these forces is zero. According to Mach, and in opposition to Newton’s conclusions, a spinning bucket of water in an empty universe would not change the water’s shape (one could not detect relative motion in an empty universe, it would be undefinable), and if the rest of the universe was set spinning while the bucket was at rest, the water surface would curve, as both the spinning of the bucket and water or the spinning of the universe are indistinguishable systems without an absolute space. Thus, for Mach, the relativity symmetry of acceleration is broken by the existence of the background universe.
@@
According to Mach, and known as Mach’s principle, the inertia of a body is not an independent and intrinsic property of matter (unlike in Newtonian mechanics and GR, where a particle in an empty universe has inertial properties), but rather the result of the action of the universe as a whole. Mach suggested that the fixed background distribution of matter in the universe must exert the inertial forces on a local accelerating body. In this way, the reference system with respect to which the universe is at rest or in uniform and rectilinear motion is a true inertial reference system. In Ernst Mach words, “I have remained to the present day the only one who insists upon referring the law of inertia to the Earth, and in the case of motions of great spatial and temporal extent to the fixed stars.[15]. Hence, inertial frames should be defined with respect to this rest frame, and local physical laws must be determined by the large-scale structure of the universe.
@@
A frame linearly accelerated relative to an inertial frame in the Minkowskian spacetime of SR is locally identical to a frame at rest in a uniform gravitational field. Einstein’s equivalence principle axiomatically assumes that inertial and gravitational masses are equal. Gravity is the only infinite range force that cannot be screened due to the existence of only positive masses, unlike in electromagnetism, where positive and negative charges result in balance and neutrality, and astrophysical and cosmological structures are therefore not governed by electromagnetism. Moreover, there always exists a reference frame in which inertial forces vanish, just as for the case of gravitational forces in free fall. These equivalences indicate that both inertial and gravitational forces are of the same nature. Following Mach, in any two-body interaction, the influence of all other matter inside their causal spacetime should be taken into account. Inertia is thus a form of gravitational induction, appearing when a body is accelerated with respect to the rest of the universe, subjected to retarded action. But Newtonian gravity and GR were developed without considering today’s acknowledgment of the mass content and size of the observable universe. Then, it seems that to describe, for example, the Earth’s motion around the Sun, only local masses and distances are required, but the Machian perspective implies that the universe’s action is already taken into account in the Newtonian laws. The only way this can be true is through the Newtonian gravitational constant, which, together with the choice of the inertial frame of absolute space, both constitute the two arbitrary choices of Newton.
@@
In Newtonian gravity, the gravitational constant $G$ is the only fundamental constant of the theory, and it is known with far less precision than any other fundamental dimensional constant in physics. The uncertainty in its measurement, together with its problematic in Quantum Field Theory in the Planck energy (which is used as a cutoff for the energy density of the vacuum and the theoretical quantum corrections estimation of the Higg’s mass, leading to the vacuum catastrophe and hierarchy problems) has led to questioning its constant nature repeatedly over the last century, see [17] for constraints. GR carries this constant (a varying $G$ would imply a violation of the strong equivalence principle) with the addition of the speed of light, which can be derived from the vacuum permittivity and permeability constants. Therefore, it is reasonable to question whether the gravitational constant is also a derived parameter.
@@
The first historically suggestion of gravity arising necessarily as a consequence of the relativity of inertia came from Hans Reissner [18]. Schrödinger soon identified in 1925 [19] that GR did not fully implement Mach’s principle, and proposed a relationship between the gravitational potential of distant masses and the speed of light: “This remarkable relationship states that the (negative) potential of all masses at the point of observation, calculated with the gravitational constant valid at that observation point, must be equal to half the square of the speed of light.”. This conclusion was reached by imposing that the kinetic energy term had an origin in a potential-like interaction (such as all other forms of energy, which have an origin in an interaction) and was relational, so that $K=mv^{2}/2$ based on absolute mass and absolute velocity can be expressed through the gravitational Newtonian potential $Gm/r$ so that $K=(Gm_{i}/r_{ij})(m_{j}v_{ij}^{2}/2c^{2})$ with the need of introducing the square of the speed of light. The term $Gm_{i}/r_{ij}c^{2}=GM_{u}/c^{2}R_{u}\sim 1$ is considered missing in the original Newtonian formulation, with $M_{u}$ being the sum of all masses with $R_{u}$, and $R_{u}$ the distance between the moving particle $m_{j}$ and $M_{u}$. Schrödinger’s formulation predicted a non-infinite observable universe, since the potential of an infinite non-expanding eternal universe would also be infinite, even though the expansion of the universe was not known at the time. This relationship was later credited to Reissner by Schrödinger himself.
@@
Schrödinger’s formulation appeared repeatedly in several works after him [20, 21, 22, 23, 24], but it was popularized by Dennis Sciama, who developed in 1953 a modified cosmological vector potential theory of inertia in which the inertial law arises as a side effect of gravity [25]. Sciama realized that in the analogous case of Maxwell’s equations applied to gravity, the rate of change of the vector potential leads to a term in the ‘gravitoelectric’ field that depends on the acceleration of an object relative to the rest of the masses of the universe. He showed that local inertial forces result from the gravitational induction of the universe, so that the dynamics of a rotating body would be affected by the large-scale distribution of the mass of the universe. Sciama postulated that “In the rest-frame of any body the total gravitational field at the body arising from all the other matter in the universe is zero”. The induced inertia decreased with $1/r$ with $r$ being the distance to the inertial body, so that the action of global matter dominates over the action of local matter, and it is not significantly modified by the acceleration with respect to local matter, leading to the illusion that inertia depends only on the body itself. The gravitational constant at any point was determined by the total potential of the distribution of matter of the universe, taking the form of $G\sim c^{2}/(\Phi_{u}+\phi)$ with $\Phi_{u}=M_{u}/R_{u}$ and local $\phi=M/r$ which reduces to $G=c^{2}/(M_{u}/R_{u})$ in most cases (when locally $\Phi_{u}>>\phi$, since only near neutron stars or black holes $\Phi_{u}\sim\phi$) with $M_{u}$ and $R_{u}$ the mass and radius of the observable universe, causally connected by the speed of light in the past light-cone of the point considered to evaluate $G$, for the equivalence of inertial and gravitational masses to hold. In this way, local phenomena were strongly coupled to global properties of the universe. Sciama did know about the Hubble expansion of the universe, but his model predicted an almost flat observable universe in which critical cosmic matter density is reached for $G\Phi_{u}/c^{2}=1$ to hold, even though the flatness of the universe was not observed at his time.
@@
By pure dimensional analysis, Reissner’s and Schrödinger’s relationship (henceforth Sciama’s relationship) is the only possible derivation of the units of measurement of the gravitational constant through a mass, a distance, and the speed of light. If Sciama’s relationship is derived from a more fundamental theory, it can be expected to also contain integers and mathematical constants, but these will be omitted for simplicity, since it does not significantly affect the resulting orders of magnitude of the comparison between $M_{u}$ and $R_{u}$.
Hans-Jürgen Treder incorporated Sciama’s mechanism by which inertia could be derived from a gravitoelectric field in a more complete inertia-free mechanics model [21] based on the Riemann potential as a velocity-dependent gravitational potential (instead of the Weber’s potential of Reissner and Schrödinger). This model implements the idea of inertia having a gravitational origin without Sciama’s anisotropic inertial mass (which is ruled out by observations) with Sciama’s $G\sim c^{2}/(\Phi_{u}+\phi)$, in agreement with Mach’s principle. In this non-relativistic linear model, the weak equivalence principle is derived and not axiomatically postulated as in GR, the weakness of the gravitational constant is explained, and inertial mass is also derived from the model.
@@
Sciama’s relationship with today’s measurements of the current cosmological model is satisfied only in terms of orders of magnitude, because the Hubble tension impedes a precise calculation. It yields the orders of magnitude of the observed gravitational constant when considering not the baryonic and dark matter content for $M_{u}$, but the total energy content (since all forms of energy must gravitate according to GR), which roughly corresponds to the critical density of the universe as measured to be almost flat. As Milgrom stated in one of his conferences, “the only system that is strongly general relativistic and in the MOND regime is the Universe at large”. However, the critical density is calculated with the assumption of a constant gravitational constant itself and considering the dark matter content, which could be reduced through an effective correction to the gravitational constant, as will be proposed in the next section. Thus, an honest calculation must be done with direct observations only by the total luminosity of the universe and its average mass-to-light ratio, which yields a value $10^{52}kg<M_{u}<10^{53}kg$, and the observable radius of the universe, which can be estimated by direct observations of the Hubble parameter (subjected to the Hubble tension), which yields around $R_{u}\sim 10^{26}m$.
@@
Few authors have attempted to develop a MOND for galaxy rotation curves based on Mach’s principle [26], or relate Milgrom’s MOND to Sciama’s relationship and Mach’s principle [27, 28, 29, 30, 31, 32, 33]. The most relevant ideas in the literature relating these topics are summarized onwards.